OpenStudy (anonymous):

How do I determine a polynomial equation whose roots include the given values: 5, -5, 1, -1

myininaya (myininaya):

f(x)=(x-5)(x-(-5))(x-1)(x-(-1)) =(x-5)(x+5)(x-1)(x+1)

OpenStudy (anonymous):

Is that the answer or the formula

OpenStudy (anonymous):

So you can just plug anything in these?

myininaya (myininaya):

no not just anything (x-the root)

OpenStudy (anonymous):

So what if I tried 1,2, 2pm sqrt{2}

OpenStudy (anonymous):

how do you do it

OpenStudy (anonymous):

I already have a midterm coming up so I need to get it

myininaya (myininaya):

f(x)=(x-1)(x-2)(x-2pmsqrt(2)) i dont know what pm means?

OpenStudy (anonymous):

the pm shows up when I use the equation sign..it's a line up and down ...it's like + with a line under it

myininaya (myininaya):

oh plus or minus

OpenStudy (anonymous):

oh, I suppose that is what the pm means then...lol

myininaya (myininaya):

f(x)=(x-1)(x-2)(x-(2+sqrt(2))(x-(2-sqrt(2))